Q1- Given a rod of a known density 6.76 g/cm3 , a mass of 392 grams,...

Q1-

Given a rod of a known density 6.76 g/cm3 , a mass of 392 grams, and a length of 0.365 meters, what is the radius of the rod in mm?

Q2-

You measured a rod to be 2.5 meters long and 5.95 mm in diameter, what is the volume of the rod in cm3?

Q3-

If you increase the length of a rod by a factor of 8 but decrease the radius by a factor of 4, by what factor does the volume of the rod change?

Q4-

You've found the following measurements for a metal rod: length = 1.82 meters, diameter = 12.02 mm, and mass = 1.46 kg. Based on these measurements, what is the density of the rod in g/cm3?
Answer #1

a)

mass = density * vol

392 = 6.76* 3.14* r^2 * 36.5

r = 0.711 cm

====

b)

V = 3.14* r^2 h

V = 3.14* (0.595 / 2)^2* 250

V =69.48 cm^3

=====

c)

since vol is directly proportional to length and sq of radius

so new vol is increased by = (1 / 4) ^2 * 8 = 0.5

=====

d)

d = 1.46* 1000 / ( 3.14* (1.202/2* 10^-2)* 182

d = 7.073 g /cm^3

======

Comment before rate in case any doubt, will reply for sure.. goodluck